OpenStudy (ttop0816):

please help! Solve secx = -1/2. no solution {120°, 240°}

OpenStudy (anonymous):

Since \(\sec x = \frac 1 {\cos x}\), what do you think you can do?

OpenStudy (ttop0816):

be a no solution?

OpenStudy (anonymous):

Yes, it is a no solution, because \(\cos x\) is restricted to \(-1 <= x <= 1\).